OpenStudy (anonymous):

Find the period and amplitude for f(x) = 2sin(3x).

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

Period is 2 right?

OpenStudy (anonymous):

Oh oops, the amplitude is 2

OpenStudy (anonymous):

@phi @sweetburger

OpenStudy (anonymous):

amplitude is 2, the absolute value of the number out front the period is \(\frac{2\pi}{b}\) which in your case is \(\frac{2\pi}{3}\)